OpenStudy (curry):

4^(1-x)=8^(2x+1) 4^(1-x)=8^(2x+1) @Mathematics

OpenStudy (anonymous):

convert everything to powers of 2:\[4^{1-x}=(2^2)^{1-x}=2^{2-2x}\]\[8^{2x+1}=(2^3)^{2x+1}=2^{6x+3}\]So now we have:\[2^{2-2x}=2^{6x+3}\iff 2-2x=6x+3\]which is easy to solve.